# The super learner for time-to-event outcomes: A tutorial

### v1

@@ 4.2.4 Extending to estimate the censoring distribution via SL
    Fit candidate learners for both $S(t|X)$ and $G(t|X)$ in a cross-validated way based on $K$ data folds, as outlined in SL steps 1-3 for $S(t|X)$ (Section 2.2). Obtain predicted probabilities from each model for each person on the time grid $V$. This gives $\hat{S}_{i,j}^{V}$, as introduced in Section 4.2.1, for $i=1,\ldots,n$, $j=1,\ldots,p$. The corresponding estimates for the censoring distribution are $\hat{G}_{i,j^{\prime}}^{V}=\left(\hat{G}_{j^{\prime}}(v|X=x_{i}),\hat{G}_{j^{\prime}}(2v|X=x_{i}),\ldots,\hat{G}_{j^{\prime}}(\tau|X=x_{i})\right)$ for $i=1,\ldots,n$, $j^{\prime}=1,\ldots,q$.
@@
    Calculate $f_{G}(i,t)=\left(1-\frac{\Delta_{i}I(\tilde{T}_{i}\leq t)}{\hat{G}(\tilde{T}_{i}|X=x_{i})}\right)$ for each person at each time point $t$ in the grid $V$ using the current estimator of the censoring distribution $G(t|X)$. Using the current values of $f^{V}_{G}(i)=\left(f_{G}(i,v),f_{G}(i,2v),\ldots,f_{G}(i,\tau)\right)$ fit a non-negative least squares regression of $f^{V}_{G}(i)$ on $\hat{S}_{i,1}^{V},\ldots,\hat{S}_{i,p}^{V}$ and scale the coefficients to sum to 1, as outlined in Section 4.2.3. This gives the current SL estimator for $S(t|X)$, $S^{SL}(t|X)=\sum_{j=1}^{p}\hat{\alpha}_{j}\hat{S}_{j}(t|X)$.
    The following steps are iterated:
@@
    Calculate $f_{S}(i,t)=\left(1-\frac{(1-\Delta_{i})I(\tilde{T}<t)}{\hat{S}(\tilde{T}|X=x_{i})}\right)$ for each person at each time point $t$ in the grid $V$ using $\hat{S}^{SL}(t|X)$ from the previous step. Fit a non-negative least squares regression of $f^{V}_{S}(i)=\left(f_{S}(i,v),f_{S}(i,2v),\ldots,f_{S}(i,\tau)\right)$ on $G_{i,1}^{V},\ldots,G_{i,q}^{V}$ and scale the coefficients to sum to 1. This gives the current SL estimator for $G(t|X)$, $G^{SL}(t|X)=\sum_{j^{\prime}=1}^{q}\hat{\beta}_{j}\hat{G}_{j^{\prime}}(t|X)$.
@@
    Calculate $f_{G}(i)$ using $\hat{G}^{SL}(t|X)$ from the previous step. Fit a non-negative least squares regression of $f^{V}_{G}(i)$ on $\hat{S}_{i,1}^{V},\ldots,\hat{S}_{i,p}^{V}$ and scale the coefficients to sum to 1, to give an updated SL estimator $S^{SL}(t|X)$.
@@
    Obtain estimates of $S(t|X)$ for all individuals for all time points in the grid $V$ based on the current $S^{SL}(t|X)$, giving $\hat{S}^{SL}(t|X=x_{i}),i=1,\ldots,n$. These estimates are now compared these with the estimates $\hat{S}^{SL}(t|X=x_{i})$ from the previous iteration. Calculate the maximum absolute difference across all $n\times|V|$ person-time combinations. If this is less than a small pre-determined value $\epsilon$ (Westling and others (2023) use $\epsilon=10^{-5}$ in the implementation in their code) then we stop. Otherwise we return to step 4 for another iteration, repeating steps 4-6 until the maximum absolute difference is less than $\epsilon$.
@@
4.3 The state learner: Munch & Gerds (2024)
@@ 4.3.1 SL steps 1-3: cross-validation
The approach of Munch and Gerds (2024) makes use of state occupation functions. In the setting of a single event of interest, at any time $t$ an individual is in one of three possible states: they have been observed to have had the event ($s=1$), they remain uncensored and have not had the event ($s=0$), or they have been censored (s=1s=-1). Let $\Lambda(t|X)$ denote the conditional cumulative hazard for the event of interest, and let $\Gamma(t|X)$ denote the conditional cumulative hazard for censoring. The conditional probabilities of being in state s=0,1,1s=0,1,-1 at time $t$ (in other words, the state occupation probabilities, the risks, or the cumulative incidences) can be written as follows:
@@
The state learner involves specifying candidate learners for the cumulative hazards of both the event and the censoring, $\Lambda(t|X)$ and $\Gamma(t|X)$. We assume $p$ candidate learners for the event ($\Lambda(t|X)$) and $q$ candidate learners for censoring ($\Gamma(t|X)$). The candidate learners could be any method for continuous time-to-event outcomes that can output a conditional cumulative hazard.
The loss function is in the form of an integrated Brier score for all three states combined, and depends on the state occupation probabilities $F(t,s,X)$ (s=1,0,1s=-1,0,1). The $F(t,s,X)$ (s=1,0,1s=-1,0,1) involve integrals and these can be estimated based on a given pair of candidate learners using a fine time grid. As for the method of Westling and others (2023) we assume a grid $V=\{v,2v,3v,\ldots,\tau-v,\tau\}$ of evenly spaced times with differences $v$. The practical steps for obtaining cross-validated estimates of $F(t,s,X)$ (s=1,0,1s=-1,0,1) are:
@@
    For individuals in validation set $D_{k}$ obtain estimates of the state occupation probabilities F(t,s,X),s=1,0,1F(t,s,X),s=-1,0,1 for each pair of models $(j,j^{\prime})$ fitted in training set $D_{-k}$:
@@ 4.3.2 SL step 4: Loss function
The estimated Brier score at time $t$ (for each $t\in V$), based on candidate learner pair $(j,j^{\prime})$ for the event and censoring cumulative hazards $\Lambda(t|X)$ and $\Gamma(t|X)$ applied to validation set $D_{k}$, is
$\hat{B}^{(j,j^{\prime},k)}(t)=\sum_{s=-1,0,1}\frac{1}{|D_{k}|}\sum_{i\in D_{k}}\left(\hat{F}_{j,j^{\prime}}(t,s,x_{i})-I(\eta_{i}(t)=s)\right)^{2},$ (20)
and the corresponding integrated Brier score is estimated based on the time grid $V$ using
@@
The estimated mean expected loss for pair of candidate learners $(j,j^{\prime})$ is $\frac{1}{K}\sum_{k=1}^{K}\widehat{IBS}^{(j,j^{\prime},k)}(\tau)$. The final state learner is given by the pair of candidate learners $(j,j^{\prime})$ with the lowest cross-validated mean expected loss estimate.
5 SL implementation in R
5.1 Discrete-time SL
@@
There is a wide choice of candidate learners available in the SuperLearner package. It includes a built in library of 40 learners. Users can also specify screening algorithms, which perform variable selection using specified methods (e.g. Lasso) before then implementing the selected candidate learners for the selected variables. There are 8 built-in screening algorithm options. The package also allows users to specify their own learners and screening algorithms. The github repository https://github.com/ecpolley/SuperLearnerExtra provides additional options.
The SuperLearner function implements the L2 loss function $L_{L2}$ (method = "method.NNLS") and the log-likelihood loss function $L_{loglik}$ (method = "method.NNloglik"), but not the IPCW loss function $L_{IPCW}$, despite $L_{IPCW}$ being recommended for use when the target estimand is a survival probability. However, previous work has found results obtained using the three loss functions to be very similar (Tanner and others, 2020). It does not directly implement the 2-stage approach to allowing flexibility in modelling the baseline hazard using a GAM. A number of other loss functions are available in the package, but these would not necessarily be suitable for discrete-time-to-event outcomes.
@@
5.2 Continuous-time SL
The iterative SL approach of Westling and others (2023) can be implemented using their survSuperLearner package, which can be installed from Github (https://github.com/tedwestling/survSuperLearner). This package accommodates a number of candidate learners: Kaplan-Meier estimator, Cox regression, parametric survival models (exponential, Weibull, log-logistic, piecewise constant hazard), generalised additive Cox model, random survival forest, and an approach that estimates a piecewise constant hazard using a separate SL for a binary outcome (using the SuperLearner package). Each of the candidate learners can be combined with different screening algorithms. The screening algorithms included are: include all covariates; include covariates selected using a Cox Lasso regression; include covariates based on separate univariable Cox regressions for each covariate (with selection based on the Wald-test p-value being below a specified threshold). For the Kaplan-Meier method no covariates are used and the same results are obtained under all screening algorithms. The code also includes a template for writing your own prediction or screening algorithm. The survSuperLearner package also allows for additional observation weights.
The state learner of Munch and Gerds (2024) can be implemented using R code from https://github.com/amnudn/joint-survival-super-learner. At the time of writing, the code works for up to two event types plus censoring, and it accommodates the following set of candidate learners for the cumulative hazards: Cox regression, the Nelson-Aalen estimator, a Cox Lasso, and a random survival forest. The statelearner code performs 5-fold cross-validation.
The continuous-time SLs use a time grid to estimate certain integrals. The grid can be specified in both survSuperLearner and statelearner. The default in survSuperLearner is to use 250 time points equally spaced between 0 and the last uncensored follow-up time (for the event survival distribution) or the last censored follow-up time minus a small number (for the censoring distribution). The statelearner defaults to 100 equally spaced time points from 0 up to the time horizon of interest.
The survSuperLearner package enable predictions to be obtained on an external data set by allowing the user to provide training and test data sets (as in SuperLearner). The statelearner uses a separate predict function that applies the selected model based on a training data set to a test data set.
@@ 6.1 Data
We illustrate the use of the SL approaches described in this paper through an application to predict conditional survival probabilities following breast cancer diagnosis using the ‘Rotterdam’ data set, which is freely available as part of the survival package in R (Therneau and others, 2024). This data set has been widely used to illustrate survival analysis methods (Royston and Altman, 2013). It includes data on individuals who underwent surgery for primary breast cancer between 1978 and 1993, and whose data were recorded in the Rotterdam Tumour Bank. Individuals were followed up for disease recurrence and death for up to a maximum of 19.3 years. The data include information on the following potential predictor variables: year of diagnosis, age, menopausal status (0: premenopausal, 1: postmenopausal), tumour size ($\leq 20$$m$, 205020-50mm, >50>50mm), differentiation grade (2 or 3), number of positive lymph nodes, progesterone receptors (continuous, fmol/l), estrogen receptors (continuous, fmol/l), chemotherapy (0: no, 1: yes), hormonal treatment (0: no, 1: yes).
@@ 6.2 Methods: implementation of the super learners
The target estimand is the conditional survival probability up to 10 years. We illustrate the use of the methods in the setting in which one data set is available for model development, and it is of interest to apply the resulting prediction algorithm in an external test data set. We divide the Rotterdam data randomly into a sample of 2087 individuals (70%) used for the development/training, with the remaining 895 individuals (30%) being used as a test set in which we will apply the SL and also evaluate its performance externally. We emphasise that using the SL involves cross-validation within the training data.
@@
To implement the methods we make use of the packages and code referenced in Section 5. Code are provided at https://github.com/ruthkeogh/superlearner_survival_tutorial. We also provide ‘by hand’ implementations (i.e. not using the packages) of the discrete-time SL method and the statelearner approach to clearly illustrate the steps. We first use a similar simple set of candidate learners across the three SL methods, though they way they are used differs slightly across methods. We then implement a richer set of candidate learners where possible using the available packages. For all SLs we used 5-fold cross-validation.
@@
For the discrete-time SL we divided time up into 10 or 100 periods of equal length to illustrate results with different levels of discretisation. When using 10 time periods we entered time $t$ as a categorical variable denoting the period, and when using 100 time periods we entered $t,t^{2},t^{3}$ as covariates. For the implementation using a simple set of learners, the candidate learners used were: a non-parametric empirical estimate including no covariates, logistic regression with and without Lasso selection of covariates, and a random forest (using the default settings of SL.ranger). We performed the analysis using the ‘L2’ and ‘loglik’ loss functions. The IPCW loss function is not accommodated in the SuperLearner package, but we illustrate its use in the by-hand implementation. For the discrete-time SL using a richer set of candidate learners we included random forests, GAMs and XGboost with different sets of tuning parameters, alongside the simple set of learners, following the set used by Tanner and others (2020) (see Supplementary Table S1).
@@
For the continuous-time SL of Westling and others (2023) we used the iterative approach in which a SL is used for both the event and censoring distributions, as is implemented in the survSuperLearner package. In the iterative process we used a value of $\epsilon=10^{-5}$ as the convergence threshold, as applied in the available code. For the implementation using a simple set of learners, the candidate learners used were the Kaplan-Meier estimator, Cox regression with and without Lasso selection of covariates, and a random survival forest (using the default settings in R, which includes to use a miniumum terminal node size of 15). We also implemented this SL using the full set of candidate learners that are built into the survSuperLearner package, which additionally include exponential, Weibull and log-logistic models, Cox regression including GAMs, and piecewise constant hazard regressions (see Supplementary Table S4).
@@
For the continuous-time SL (‘state learner’) of Munch and Gerds (2024) we used the same set of candidate learners for the event and censoring distributions. We used the Kaplan-Meier estimator, Cox regression with and without Lasso selection of covariates, and random survival forest, which we understand to be the full set of built-in learners in statelearner at the time of writing. The continuous-time SLs use a time grid to estimate certain integrals, and we used the defaults for this. In the by-hand implementation we evaluated the integrals using all observed event times.
@@
Using the continuous-time data we also applied each of the simple candidate learners (Cox regression, Cox regression with Lasso selection of covariates, random survival forest) separately, for comparison with the SLs.

### v2

@@ 4.2.4 Extending to estimate the censoring distribution via super learner
    Fit candidate learners for both $S(t|X)$ and $G(t|X)$ in a cross-validated way based on $K$ data folds, as outlined in SL steps 1-3 for $S(t|X)$ (Section 2.2). Obtain predicted probabilities from each model for each person on the time grid $V$. This gives $\hat{S}_{i,j}^{V}$, as introduced in Section 4.2.1, for $i=1,\ldots,n$; $j=1,\ldots,p$. The corresponding estimates for the censoring distribution are $\hat{G}_{i,j^{\prime}}^{V}=\left(\hat{G}_{j^{\prime}}(v|X=x_{i}),\hat{G}_{j^{\prime}}(2v|X=x_{i}),\ldots,\hat{G}_{j^{\prime}}(\tau|X=x_{i})\right)$ for $i=1,\ldots,n$; $j^{\prime}=1,\ldots,q$.
@@
    Calculate $\hat{f}_{G}(i,t)=\left(1-\frac{\Delta_{i}I(\tilde{T}_{i}\leq t)}{\hat{G}(\tilde{T}_{i}|X=x_{i})}\right)$ for each person at each time point $t$ in the grid $V$ using the current estimator of the censoring distribution $G(t|X)$. Using the current values of $\hat{f}^{V}_{G}(i)=\left(\hat{f}_{G}(i,v),\hat{f}_{G}(i,2v),\ldots,\hat{f}_{G}(i,\tau)\right)$ fit a non-negative least squares regression of $\hat{f}^{V}_{G}(i)$ on $\hat{S}_{i,1}^{V},\ldots,\hat{S}_{i,p}^{V}$ and scale the coefficients to sum to 1, as outlined in Section 4.2.3. This gives the current SL estimator for $S(t|X)$, $S^{SL}(t|X)=\sum_{j=1}^{p}\hat{\alpha}_{j}\hat{S}_{j}(t|X)$.
    The following steps are then iterated:
@@
    Calculate $\hat{f}_{S}(i,t)=\left(1-\frac{(1-\Delta_{i})I(\tilde{T}<t)}{\hat{S}(\tilde{T}|X=x_{i})}\right)$ for each person at each time point $t$ in the grid $V$ using $\hat{S}^{SL}(t|X)$ from the previous step. Fit a non-negative least squares regression of $\hat{f}^{V}_{S}(i)=\left(\hat{f}_{S}(i,v),\hat{f}_{S}(i,2v),\ldots,\hat{f}_{S}(i,\tau)\right)$ on $G_{i,1}^{V},\ldots,G_{i,q}^{V}$ and scale the coefficients to sum to 1. This gives the current SL estimator for $G(t|X)$, $G^{SL}(t|X)=\sum_{j^{\prime}=1}^{q}\hat{\beta}_{j}\hat{G}_{j^{\prime}}(t|X)$.
@@
    Calculate $\hat{f}_{G}(i)$ using $\hat{G}^{SL}(t|X)$ from the previous step. Fit a non-negative least squares regression of $\hat{f}^{V}_{G}(i)$ on $\hat{S}_{i,1}^{V},\ldots,\hat{S}_{i,p}^{V}$ and scale the coefficients to sum to 1, to give an updated SL estimator $S^{SL}(t|X)$.
@@
    Obtain estimates of $S(t|X)$ for all individuals for all time points in the grid $V$ based on the current $S^{SL}(t|X)$, giving $\hat{S}^{SL}(t|X=x_{i}),i=1,\ldots,n$. These estimates are now compared these with the estimates $\hat{S}^{SL}(t|X=x_{i})$ from the previous iteration. Calculate the maximum absolute difference across all $n\times|V|$ person-time combinations, where $|V|$ is the number of time in the grid $V$. If the maximum absolute difference is less than a small pre-determined value $\epsilon$ then we stop. Westling and others (2023) use $\epsilon=10^{-5}$ in the implementation in their code. Otherwise we return to step 4 for another iteration, repeating steps 4-6 until the maximum absolute difference is less than $\epsilon$.
@@
4.3 The joint survival super learner: Munch & Gerds (2025)
@@ 4.3.1 SL steps 1-3: cross-validation
The joint survival SL approach of Munch and Gerds (2025) makes use of state occupation functions. In our setting of a single event of interest, at any time $t$ an individual in $D$ can be considered to be in one of three possible states denoted $s$: (i) they remain uncensored and have not had the event ($s=0$), (ii) they have been observed to have had the event ($s=1$), or (iii) they have been censored (s=1s=-1). Let $\Lambda(t|X)$ denote the conditional cumulative hazard for the event of interest, and let $\Gamma(t|X)$ denote the conditional cumulative hazard for censoring. The conditional probabilities of being in state s=0,1,1s=0,1,-1 at time $t$ (in other words, the state occupation probabilities, the risks, or the cumulative incidences) can be written as follows:
@@
The joint survival SL involves specifying candidate learners for the cumulative hazards of both the event and the censoring, $\Lambda(t|X)$ and $\Gamma(t|X)$. We assume $p$ candidate learners for the event ($\Lambda(t|X)$) and $q$ candidate learners for censoring ($\Gamma(t|X)$). The candidate learners could be any method for continuous time-to-event outcomes that can output a conditional cumulative hazard. Munch and Gerds (2025) consider Cox regression, the Nelson-Aalen estimator, a Cox Lasso, and a random survival forest.
As we will see in Section 4.3.2, the loss function is in the form of an integrated Brier score for all three states combined, and depends on the state occupation probabilities $F(t,s,X)$ (s=1,0,1s=-1,0,1). To estimate the mean expected loss we need to obtain cross-validated estimates of $F(t,s,X=x_{i})$ (s=1,0,1s=-1,0,1) for each individual in $D$. The $F(t,s,X)$ (s=1,0,1s=-1,0,1) involve integrals and these can be estimated based on a given pair of candidate learners (for $\Lambda(t|X)$ and $\Gamma(t|X)$) using a fine time grid. As for the method of Westling and others (2023) we assume a grid $V=\{v,2v,3v,\ldots,\tau-v,\tau\}$ of evenly spaced times with differences $v$. The practical steps for obtaining cross-validated estimates of $F(t,s,X)$ (s=1,0,1s=-1,0,1) are:
@@
    For individuals in validation set $D_{k}$ obtain estimates of the state occupation probabilities F(t,s,X),s=1,0,1F(t,s,X),s=-1,0,1 for each pair of candidate learners $(j,j^{\prime})$ fitted in training set $D_{-k}$:
@@ 4.3.2 SL step 4: Loss function
The estimated Brier score at time $t$ based on candidate learner pair $(j,j^{\prime})$ for the event and censoring cumulative hazards $\Lambda(t|X)$ and $\Gamma(t|X)$ applied to validation set $D_{k}$, is
$\hat{B}^{(j,j^{\prime},k)}(t)=\sum_{s=-1,0,1}\frac{1}{|D_{k}|}\sum_{i\in D_{k}}\left(\hat{F}_{j,j^{\prime}}(t,s,x_{i})-I(\eta_{i}(t)=s)\right)^{2}.$ (20)
This can be calculated for each $t$ in the time grid $V$ and the corresponding integrated Brier score is estimated using
@@
The estimated mean expected loss for the pair of candidate learners $(j,j^{\prime})$ is $\frac{1}{K}\sum_{k=1}^{K}\widehat{IBS}^{(j,j^{\prime},k)}(\tau)$.
4.3.3 The non-ensemble super learner
The approach of Munch and Gerds (2025) is designed to derive a non-ensemble SL. The final joint survival SL is given by the pair of candidate learners $(j,j^{\prime})$ with the lowest cross-validated mean expected loss estimate based on (21). In our setting where the target estimand is $S(\tau|X=x)=\Pr(T>\tau|X=x)$ the SL estimator is given by $S^{SL}(\tau|X)=\exp(-\hat{\Lambda}_{j}(\tau|X))$, where $j$ is the selected candidate learner for $\Lambda(t|X)$ and the estimates $\hat{\Lambda}_{j}(\tau|X)$ are obtained based on fitting candidate learner $j$ to the full data $D$. The selected candidate learner for the censoring distribution is not used in the estimator $S^{SL}(\tau|X)$, but the selection of candidate learner $j$ for $\Lambda(t|X)$ was based on considering pairs of candidate learners for $\Lambda(t|X)$ and $\Gamma(t|X)$.
5 Super learner implementation in R
5.1 Discrete-time super learner
@@
There is a wide choice of candidate learners available in the SuperLearner package. It includes a built in library of 40 learners. Users can also specify screening algorithms, which perform variable selection using specified methods (e.g. Lasso) before then implementing the selected candidate learners for the selected variables. The SuperLearner package includes 8 built-in screening algorithm options. The package also allows users to specify their own learners and screening algorithms. The github repository https://github.com/ecpolley/SuperLearnerExtra provides additional options.
Three loss functions for use in the discrete-time SL approach were discussed in Section 3.3. The SuperLearner function implements the L2 loss function $L_{L2}$ (using method = "method.NNLS") and the log-likelihood loss function $L_{loglik}$ (using method = "method.NNloglik"), but not the IPCW loss function $L_{IPCW}$, despite $L_{IPCW}$ being recommended for use when the target estimand is a survival probability (Polley and van der Laan, 2011). However, previous work has found results obtained using the three loss functions to be very similar (Tanner and others, 2020). A number of other loss functions are available in the package, but these would not necessarily be suitable for discrete-time-to-event outcomes. The SuperLearner package does not directly implement the 2-stage approach to allowing flexibility in modelling the baseline hazard using a GAM.
@@
5.2 Continuous-time super learner
The iterative SL approach of Westling and others (2023) outlined in Section 4.2.4 can be implemented using their survSuperLearner package, which can be installed from Github (https://github.com/tedwestling/survSuperLearner). This package accommodates a number of candidate learners: Kaplan-Meier estimator, Cox regression, parametric survival models (exponential, Weibull, log-logistic, piecewise constant hazard), generalised additive Cox model, random survival forest, and an approach that estimates a piecewise constant hazard using a separate SL for a binary outcome (implemented using the SuperLearner package). The number of folds to be used in the cross-validation can be specified as an option in the function. Each of the candidate learners can be combined with different screening algorithms. The screening algorithms included are: include all covariates; include covariates selected using a Cox Lasso regression; include covariates based on separate univariable Cox regressions for each covariate (with selection based on the Wald-test p-value being below a specified threshold). For the Kaplan-Meier method no covariates are used and the same results are obtained under all screening algorithms. The code also includes a template for writing your own prediction or screening algorithm. The survSuperLearner package also allows for additional observation weights.
The joint survival SL of Munch and Gerds (2025) can be implemented using R code from https://github.com/amnudn/joint-survival-super-learner (jossl). The code currently accommodates up to two event types plus censoring, and it accommodates the following set of candidate learners for the cumulative hazards: Cox regression, the Nelson-Aalen estimator, a Cox Lasso, and a random survival forest. The jossl code performs 5-fold cross-validation.
The continuous-time SLs use a time grid to estimate certain integrals (denoted $V$ in Section 4). The grid can be specified in both survSuperLearner and jossl. The default in survSuperLearner is to use 250 time points equally spaced between 0 and the last uncensored follow-up time (for the event survival distribution) or the last censored follow-up time minus a small number (for the censoring distribution). In jossl the default is 100 equally spaced time points from 0 up to the time horizon of interest.
The survSuperLearner package enables predictions to be obtained on an external data set by allowing the user to provide training and test data sets, as in SuperLearner.
@@ 6.1 Data
We illustrate the use of the three SL approaches described in this paper through an application to predict conditional survival probabilities following breast cancer diagnosis using the ‘Rotterdam’ data set, which is freely available as part of the survival package in R (Therneau and others, 2024). This data set has been widely used to illustrate survival analysis methods (Royston and Altman, 2013). It includes data on individuals who underwent surgery for primary breast cancer between 1978 and 1993, and whose data were recorded in the Rotterdam Tumour Bank. Individuals were followed up for disease recurrence and death for up to a maximum of 19.3 years. The data include information on the following potential predictor variables: year of diagnosis, age, menopausal status (0: premenopausal, 1: postmenopausal), tumour size ($\leq 20$$m$, 205020-50mm, >50>50mm), differentiation grade (2 or 3), number of positive lymph nodes, progesterone receptors (continuous, fmol/l), estrogen receptors (continuous, fmol/l), chemotherapy (0: no, 1: yes), hormonal treatment (0: no, 1: yes).
@@ 6.2 Methods: implementation of the super learners
The target estimand is the conditional survival probability up to 10 years. We illustrate the use of the methods in the setting in which a training data set is available for model development, and it is of interest to apply the resulting prediction algorithm in an external test data set. We divide the Rotterdam data randomly into a sample of 2087 individuals (70%) used for the development/training, with the remaining 895 individuals (30%) being used as a test set in which we will apply the SL and also evaluate its performance externally. We emphasise that using the SL involves cross-validation within the training data.
@@
To implement the methods we make use of the packages and code referenced in Section 5. Code for reproducing the analyses and results is provided at https://github.com/ruthkeogh/superlearner_survival_tutorial. We also provide ‘by hand’ implementations (i.e. not using the packages) of the discrete-time SL method and the joint survival SL approach to clearly illustrate the steps. We first use a similar simple set of candidate learners across the three SL methods, though they way they are used differs slightly across methods. We then implement a richer set of candidate learners where possible using the available packages. For all SLs we used 5-fold cross-validation.
@@
For the discrete-time SL of Polley and van der Laan (2011) we divide time up into 10 or 100 periods of equal length to illustrate results with different levels of discretisation. When using 10 time periods we enter time $t$ as a categorical variable for each period, and when using 100 time periods we enter $t,t^{2},t^{3}$ as continuous variables. For the implementation using a simple set of learners, the candidate learners used are: a non-parametric empirical estimate including no covariates, logistic regression with and without Lasso selection of covariates, and a random forest (using the default settings of SL.ranger). We perform the analysis using the $L2$ and loglikloglik loss functions. The IPCW loss function is not accommodated in the SuperLearner package, but we illustrate its use in our by-hand implementation. For the discrete-time SL using a richer set of candidate learners we include random forests, GAMs and XGboost with different sets of tuning parameters, alongside the simple set of learners, following the set used by Tanner and others (2020) (see Supplementary Table S3).
@@
For the continuous-time SL of Westling and others (2023) we use the iterative approach in which a SL is used for both the event and censoring distributions, as is implemented in the survSuperLearner package. In the iterative process we use a value of $\epsilon=10^{-5}$ as the convergence threshold, as applied in the available code. For the implementation using a simple set of learners, the candidate learners used are the Kaplan-Meier estimator, Cox regression with and without Lasso selection of covariates, and a random survival forest (using the default settings in R, which includes to use a miniumum terminal node size of 15). We also implement this SL using the full set of candidate learners that are built into the survSuperLearner package, which additionally include exponential, Weibull and log-logistic models, Cox regression including GAMs, and piecewise constant hazard regressions (see Supplementary Table S4).
@@
For the continuous-time SL (‘joint survival SL’) of Munch and Gerds (2025) we use the same set of candidate learners for the event and censoring distributions. These are the Nelson-Aalen estimator, Cox regression with and without Lasso selection of covariates, and random survival forest. This gives 16 possible pairs of candidate learners for the event and censoring distributions.
The continuous-time SLs use a time grid to estimate certain integrals, and we use the defaults for this. In the by-hand implementation we evaluate the integrals using a time grid defined by all observed event times.
@@
Using the continuous-time data we also apply each of the simple candidate learners (Cox regression, Cox regression with Lasso selection of covariates, random survival forest) separately, for comparison with the SLs.